# AVTENet: Audio-Visual Transformer-based Ensemble Network Exploiting Multiple Experts for Video Deepfake Detection

### v1

@@ IV-B Hyperparameters
We use the Adam optimizer to train each classifier. AN (based on AST) is trained with a learning rate of 0.00001, VN (based on ViViT) is trained with a learning rate of 0.0001 and AVN (with AV-HuBERT) is trained with a learning rate of 0.002. Similarly, AVN (with AST&ViViT) is trained with a learning rate of 0.0001. The linear layer in AVTENet for combining AST-based AN, ViViT-based VN, and AV-HuBERT-based AVN is trained with a learning rate of 0.002, while the linear layer in AVTENet for combining AST-based AN, ViViT-based VN and AST&ViViT-based AVN is trained with a learning rate of 0.002.
@@ IV-C Evaluation Metrics
We use precision, recall, F1-score, and accuracy to evaluate the performance of individual uni-modal/multi-modal classifiers and our proposed multi-modal AVTENet classifier. For all metrics, a higher value indicates better performance. For a given test input, a binary classifier produces a binary prediction output, either 1 (for positive) or 0 (for negative). Compared to the ground-truth label, the prediction can be true positive (TP: correct positive prediction), false positive (FP: incorrect positive prediction), true negative (TN: correct negative prediction), and false negative (FN: incorrect negative prediction).
Accuracy is the primary metric. It calculates the proportion of correct predictions out of total predictions:
$\mathrm{Accuracy}={\frac{TP+TN}{TP+TN+FP+FN}}.$ (6)
The precision score determines the capability of the model to correctly identify positive samples. It is a measure of the proportion of true positive predictions to the total number of positive predictions:
$\mathrm{Precision}={\frac{TP}{TP+FP}}.$ (7)
The recall score, also known as sensitivity or true positive rate, measures how much a model captures positive instances. It is an estimate of the proportion of true positive predictions to the actual positive instances in the dataset:
$\mathrm{Recall}={\frac{TP}{TP+FN}}.$ (8)
The F1-score is the harmonic mean of precision and recall, which combines them in one metric. It is calculated by taking into account false positives and false negatives:
[…]
@@ IV-D Experimental Results
This section provides detailed experimental results and their analysis. Testset-I and Test-II are the main evaluation test sets of the FakeAVCeleb dataset.
IV-D1 Video-only Detection Results
@@
Testset-I Real 0.87 0.87 0.87 0.87
Fake 0.87 0.87 0.87
Testset-II Real 0.76 0.87 0.81 0.80
Fake 0.85 0.73 0.78
The detailed results of the VN classifier on eight different test sets are shown in Table IV. We observe that the VN classifier has better performance on the test sets containing visual manipulations because the VN classifier utilizes visual features to detect forgeries in videos. The accuracy on the faceswap, faceswap-wav2lip, fsgan, fsgan-wav2lip, and wav2lip test sets is all above 0.9, except for RTVC, which achieves an accuracy of 0.5. Table I evidently shows that, except for RTVC, all these test sets contain visual manipulations, making VN perform better on these test sets. The poor performance on RTVC is due to the lack of visual manipulations. Here, it is important to highlight that besides visual manipulations, both faceswap-wav2lip and fsgan-wav2lip also contain acoustic manipulation, but the VN classifier only detects visual manipulation in these instances. Furthermore, the accuracy of the two main test sets (0.87 for Test set-I and 0.80 for Testset-II) is relatively low compared to other test sets. This is because, for both test sets, the fake class contains video instances with or without video manipulation. The conclusion drawn from this experiment is that the VN classifier performs well when the video is visually manipulated but fails completely when the video only contains acoustic manipulation.
@@
Testset-I Real 0.71 1.00 0.83 0.80
Fake 1.00 0.60 0.75
Testset-II Real 0.76 1.00 0.86 0.84
Fake 1.00 0.69 0.81
IV-D2 Audio-only Detection Results
The AN classifier is a uni-modal method that uses acoustic features to detect whether a video is real or fake. The detailed results of AN on eight different test sets are shown in Table V. We observe a similar pattern to the results of the VN classifier in Table IV. The detection accuracy is almost perfect on the faceswap-wav2lip, fsgan-wav2lip, and RTVC test sets (all containing acoustic manipulation in the fake class). But AN fails completely on the faceswap and fsgan test sets, where fake instances are only visually manipulated while the audio tracks are all real. Furthermore, the detection accuracies of Testset-I and Testset-II (0.80 and 0.84, respectively) are lower than the strong performance of the other test sets. This is because, for both test sets, the fake class contains video instances with or without acoustic manipulation. The results of this experiment show that the AN classifier can detect acoustically manipulated videos well, but cannot detect fake videos containing only visual manipulations.
@@ IV-D2 Audio-only Detection Results
Testset-I Real 0.92 0.96 0.94 0.94
Fake 0.96 0.91 0.93
Testset-II Real 1.00 0.96 0.98 0.98
Fake 0.96 1.00 0.98
IV-D3 Audio-Visual Detection Results
Compared to the VN and AN classifiers, which can only detect visual and acoustic manipulations in fake videos, respectively, the AVN classifier utilizes both acoustic and visual information to detect forgeries in videos. Table VI shows the detailed results of the AV-HuBERT-based AVN classifier on eight different test sets. Several observations can be drawn from the table. First, the AVN classifier achieves good detection performance on all test sets except the faceswap test set. One possible reason is that the faceswap test set only involves visual manipulation, and the AV-HuBERT feature extractor only extracts visual features from lip images, thereby losing information outside the lip region. Second, although the AV-HuBERT-based AVN classifier performs well on the test sets containing manipulation in a single modality, the performance may be slightly lower than audio-only and video-only classifiers focusing on one modality. For example, the AVN classifier achieves an accuracy of 0.96 on RTVC (containing only acoustic manipulation), but the AN classifier achieves an accuracy of 0.99 on the same test (see Table V). Third, the AV-HuBERT-based AVN classifier achieves accuracy of 0.94 and 0.98 on Testset-I and Testset-II, whose fake instances contain various manipulations in acoustic and visual streams, outperforming the VN classifier (cf. 0.87 and 0.80 in Table IV) and the AN classifier (cf. 0.80 and 0.84 in Table V). Therefore, the insight gleaned from this experiment is that the AVN classifier successfully detects forgeries in videos by taking into account both acoustic and visual information, generally outperforming VN and AN classifiers that focus on uni-modal visual or acoustic manipulation.
@@ IV-D3 Audio-Visual Detection Results
TABLE VIII: Detection Results of $AVTENet_{asf}$ (Average Score fusion).
Test-Set type Class Precision Recall F1-Score Accuracy
@@
Testset-I Real 0.92 1.00 0.96 0.96
Fake 1.00 0.91 0.96
Testset-II Real 1.00 1.00 1.00 1.00
Fake 1.00 1.00 1.00
IV-D4 Detection Results of AVTENet
$\mathbf{AVTENet_{mv}}$: Table VII shows the detailed results of $AVTENet_{mv}$ on eight different test sets. $AVTENet_{mv}$ achieves perfect accuracy of 1.00 on faceswap-wav2lip, fsgan-wav2lip, and Testset-II and high accuracy of 0.99, 0.98, and 0.96 on wav2lip, RTVC, and Testset-I. This result confirms the advantages of integrating VN, AN, and AVN networks in fully considering visual and acoustic information in fake video detection. However, $AVTENet_{mv}$ achieves relatively low performances on faceswap and fsgan. As explained before, this may be because the faceswap and fsgan test sets only involve visual manipulations, while the AV-HuBERT feature extractor ignores visual information outside the lip region. Comparing Table VII with Tables IV, V, and VI, $AVTENet_{mv}$ outperforms all its component networks in all test conditions except for VN on faceswap and AN on RTVC. This result shows that the ensemble $AVTENet_{mv}$ comes with some tradeoffs in fake detection compared to modality-specific classifiers operating on the corresponding single-modality manipulation. Finally, it is worth mentioning again that $AVTENet_{mv}$ outperforms all its component networks on the two main tests, Testset-I and Testset-II.
$\mathbf{AVTENet_{asf}}$: Table VIII shows the detailed results of $AVTENet_{asf}$ on eight different test sets. Comparing Table VIII and Table VII, we can see that $AVTENet_{asf}$ performs almost the same as $AVTENet_{mv}$, except that the former achieves a slightly lower accuracy than the latter on the faceswap test set (0.86 vs 0.87).
@@ IV-D4 Detection Results of AVTENet
Testset-I Real 0.96 0.91 0.93 0.94
Fake 0.92 0.96 0.94
Testset-II Real 1.00 0.91 0.96 0.96
Fake 0.92 1.00 0.96
@@
Testset-I Real 0.96 0.97 0.96 0.96
Fake 0.97 0.96 0.96
Testset-II Real 1.00 0.97 0.99 0.99
Fake 0.97 1.00 0.99
$\mathbf{AVTENet_{sf}}$: Table IX shows the detailed results of $AVTENet_{sf}$ on eight different test sets. Comparing Table IX with Tables VII and VIII, we can see that $AVTENet_{sf}$ performs more stably than $AVTENet_{mv}$ and $AVTENet_{asf}$ across 8 test sets. The stability may be attributed to the additional linear layer taking as input the scores output by the VN, AN, and AVN networks, as it learns to balance the contributions of these three component networks. However, although $AVTENet_{sf}$ outperforms $AVTENet_{mv}$ and $AVTENet_{asf}$ on the faceswap test set (0.92 vs 0.87 and 0.86), it performs worse than $AVTENet_{mv}$ and $AVTENet_{asf}$ on all the other test sets.
$\mathbf{AVTENet_{ff}}$: Table X shows the detailed results of $AVTENet_{ff}$ on eight different test sets. We can see that $AVTENet_{ff}$ performs best among four ensemble models on faceswap, fsgan, and Testset-II, and achieves near-perfect performance on faceswap-wav2lip, fsgan-wav2lip, wav2lip, and Testset-II. While the other ensemble models may excel in performance on specific test sets, $AVTENet_{ff}$ maintains stable and reliable performance on all eight test sets. The good performance may be attributed to the additional linear layer that takes as input the embeddings of the penultimate layers of the VN, AN, and AVN networks, as it learns to balance the contributions of these three component networks. In summary, feature fusion is the best fusion strategy in this work.
Comparison of Our Models
Fig. 6 depicts the accuracy of the above seven models evaluated on two main test sets, Tesetset-I and Testset-II. The video-only model (VN) is good at identifying visual manipulations in videos. Likewise, the audio-only model (AN) is good at identifying manipulated acoustic content in videos. But they only focus on a single modality. In contrast to VN and AN, the audio-visual model (AVN) and all variants of AVTENet are able to identify acoustic and visual manipulations in videos. Although $AVTENet_{mv}$ and $AVTENet_{asf}$ have higher accuracy than $AVTENet_{ff}$ on Testset-II, the latter shows higher stability in detecting various manipulation types in videos.
@@
TABLE XII: Performance comparison of our Ensemble model and several existing uni-modal, multi-modal, fusion, and ensemble models. DFD in the first column refers to the deepfake detection method. “V”, “A” and “AV” stand for visual, audio and audio-visual modalities, respectively.
@@
Unimodal[56] VGG16 V Real 0.6935 0.8966 0.7821 0.8103
Fake 0.8724 0.7750 0.8208
Unimodal[56] Xception A Real 0.8750 0.6087 0.7179 0.7626
Fake 0.7033 0.9143 0.7950
@@
Ensemble (Soft-Voting)[56] VGG16 AV Real 0.6935 0.8966 0.7821 0.7804
Fake 0.8948 0.6894 0.7788
Ensemble (Hard-Voting)[56] VGG16 AV Real 0.6935 0.8966 0.7821 0.7804
Fake 0.8948 0.6894 0.7788
Multimodal-1[56] Multimodal-1 AV Real 0.000 0.000 0.000 0.5000
Fake 0.496 1.000 0.663
Multimodal-2[56] Multimodal-2 AV Real 0.710 0.587 0.643 0.674
Fake 0.648 0.760 0.700
Multimodal-3[56] CDCN AV Real 0.500 0.068 0.120 0.515
Fake 0.500 0.940 0.651
@@
Modality Mixing [58] Multimodaltrace AV Real - - - 0.929
@@
Ensemble[59] AVFakeNet AV Real - - - 0.934
@@
Fusion[60] AVoiD-DF AV Real - - - 0.837
@@
Fusion[61] MIS-AVioDD AV Real - - - 0.962
@@
IV-D5 Comparison of AVTENet with two different AVNs

### v2

@@ IV-B Hyperparameters
We use the Adam optimizer to train each classifier. AN (based on AST) is trained with a learning rate of 0.00001, VN (based on ViViT) is trained with a learning rate of 0.0001, and AVN (with AV-HuBERT) is trained with a learning rate of 0.002. Similarly, AVN (with AST&ViViT) is trained with a learning rate of 0.0001. In AVTENet, the linear layers for combining AST-based AN, ViViT-based VN, and AVN (both AVHuBERT-based and AST&ViViT-based) are trained with the learning rate of 0.002.
@@ IV-C Evaluation Metrics
We use accuracy, precision, recall, and F1-score to evaluate the performance of individual unimodal/multimodal classifiers and our AVTENet classifier. They are calculated as follows:
${Accuracy}={\frac{TP+TN}{TP+TN+FP+FN}},$ (6)
${Precision}={\frac{TP}{TP+FP}},$ (7)
${Recall}={\frac{TP}{TP+FN}},$ (8)
${F1-score}={\frac{2\times Precision\times Recall}{Precision+Recall}},$ (9)
where ${TP}$, ${TN}$, ${FP}$, and ${FN}$ denote True Positive (fake videos correctly detected as fake), True Negative (real videos correctly detected as real), False Positive (real videos incorrectly detected as fake), and False Negative (fake videos incorrectly detected as real), respectively. For all metrics, a higher value indicates better performance.
@@ IV-D Experimental Results
This section provides detailed experimental results and analysis. Testset-I and Test-II are the main evaluation test sets of the FakeAVCeleb dataset.
IV-D1 Results of the Video-only Network
@@
Testset-I Real 0.87 0.87 0.87 0.87
Fake 0.87 0.87 0.87
Testset-II Real 0.76 0.87 0.81 0.80
Fake 0.85 0.73 0.78
@@
The results of the VN classifier on eight test sets are shown in Table IV. We observe that the VN classifier performs better on the test sets with visual manipulations because it utilizes visual features to detect forgeries in videos. The accuracy on the faceswap, faceswap-wav2lip, fsgan, fsgan-wav2lip, and wav2lip test sets are all above 0.90, except for RTVC, which achieves an accuracy of 0.50. Table I shows that, except for RTVC, all these test sets contain visual manipulations, making VN perform better on these test sets. The poor performance on RTVC is due to the lack of visual manipulations. Notably, both faceswap-wav2lip and fsgan-wav2lip include acoustic manipulations as well, but the VN classifier only detects visual manipulations in these cases. Furthermore, the accuracy of the two main test sets (0.87 for Test set-I and 0.80 for Testset-II) is relatively low compared to other test sets. This is because, for both test sets, the fake class contains videos with or without visual manipulation. In conclusion, the VN classifier performs better when the video is visually manipulated, but fails when the video only contains acoustic manipulation.
IV-D2 Results of the Audio-only Network
The AN classifier is a unimodal method that uses acoustic features to detect whether a video is real or fake. The results of AN on eight test sets are shown in Table V. We observe a similar pattern to the results of the VN classifier in Table IV. The detection accuracy is almost perfect on the faceswap-wav2lip, fsgan-wav2lip, and RTVC test sets (all containing acoustic manipulation in the fake class). However, AN fails on the faceswap and fsgan test sets, where fake instances are only visually manipulated while the audio tracks are all real. Furthermore, the detection accuracies of Testset-I and Testset-II (0.80 and 0.84, respectively) are lower than the strong performance of the other test sets. This is because, for both test sets, the fake class contains videos with or without acoustic manipulation. The results how that the AN classifier can effectively detect acoustically manipulated videos, but cannot detect fake videos containing only visual manipulations.
@@ IV-D2 Results of the Audio-only Network
Testset-I Real 0.71 1.00 0.83 0.80
Fake 1.00 0.60 0.75
Testset-II Real 0.76 1.00 0.86 0.84
Fake 1.00 0.69 0.81
@@
Testset-I Real 0.92 0.96 0.94 0.94
Fake 0.96 0.91 0.93
Testset-II Real 1.00 0.96 0.98 0.98
Fake 0.96 1.00 0.98
@@
Testset-I Real 0.92 1.00 0.96 0.96
Fake 1.00 0.91 0.96
Testset-II Real 1.00 1.00 1.00 1.00
Fake 1.00 1.00 1.00
@@
IV-D3 Results of the Audio-Visual Network
Compared to the VN and AN classifiers, which can only detect visual and acoustic manipulations in fake videos, respectively, the AVN classifier utilizes both acoustic and visual information to detect forgeries in videos. Table VI shows the results of the AV-HuBERT-based AVN classifier on eight test sets. Several observations can be drawn from the table. First, the AVN classifier achieves good detection performance on all test sets except the faceswap test set. One possible reason is that the faceswap test set only involves visual manipulation, and the AV-HuBERT feature extractor only extracts visual features from lip images, thereby losing information outside the lip region. Second, although the AV-HuBERT-based AVN classifier performs well on the test sets containing manipulation in a single modality . However, its performance may be slightly lower than audio-only and video-only classifiers focusing on one modality. For instance, the AVN classifier achieves an accuracy of 0.96 on RTVC (fake samples containing only acoustic manipulation), but the AN classifier achieves an accuracy of 0.99 on the same test set (see Table V). This is because the AN classifier only focuses on the acoustic modality, while the AVN classifier will inevitably be interfered with by the visual feature extracted from the real visual modality when detecting acoustic manipulation. Third, the AV-HuBERT-based AVN classifier achieves accuracy of 0.94 and 0.98 on Testset-I and Testset-II, whose fake instances contain various manipulations in acoustic and visual streams, outperforming the VN classifier (cf. 0.87 and 0.80 in Table IV) and the AN classifier (cf. 0.80 and 0.84 in Table V). The results reveal that the AVN classifier successfully detects video forgeries by considering both acoustic and visual information. It generally outperforms VN and AN classifiers, which focus on only visual or acoustic manipulation.
IV-D4 Results of AVTENet
TABLE VIII: Detection Results of $AVTENet_{asf}$ (Average
Score fusion).
Test-Set type Class Precision Recall F1-Score Accuracy
@@ IV-D4 Results of AVTENet
Testset-I Real 0.96 0.91 0.93 0.94
Fake 0.92 0.96 0.94
Testset-II Real 1.00 0.91 0.96 0.96
Fake 0.92 1.00 0.96
@@
$\mathbf{AVTENet_{mv}}$: Table VII shows the results of $AVTENet_{mv}$ on eight different test sets. $AVTENet_{mv}$ achieves perfect accuracy of 1.00 on faceswap-wav2lip, fsgan-wav2lip, and Testset-II and high accuracy of 0.99, 0.98, and 0.96 on wav2lip, RTVC, and Testset-I. This result confirms the advantages of integrating VN, AN, and AVN networks in fully considering visual and acoustic information in fake video detection. However, $AVTENet_{mv}$ achieves relatively low performance on faceswap and fsgan. As explained before, this may be because the faceswap and fsgan test sets only involve visual manipulations, while the AV-HuBERT feature extractor ignores visual information outside the lip region. Comparing Table VII with Tables IV, V, and VI, $AVTENet_{mv}$ outperforms all its component networks in all test conditions except for VN on faceswap and AN on RTVC. This result shows that the ensemble $AVTENet_{mv}$ comes with some tradeoffs in fake detection compared to modality-specific classifiers operating on the corresponding unimodal manipulation.
$\mathbf{AVTENet_{asf}}$: Table VIII shows the results of $AVTENet_{asf}$ on eight different test sets. Comparing Table VIII and Table VII, we can see that $AVTENet_{asf}$ performs almost the same as $AVTENet_{mv}$, except that the former achieves a slightly lower accuracy than the latter on the faceswap test set (0.86 vs 0.87).
$\mathbf{AVTENet_{sf}}$: Table IX shows the results of $AVTENet_{sf}$ on eight different test sets. Comparing Table IX with Tables VII and VIII, we can see that $AVTENet_{sf}$ performs more stably than $AVTENet_{mv}$ and $AVTENet_{asf}$ across test sets. The stability may be attributed to the additional linear layer taking as input the output scores of the VN, AN, and AVN networks, as it learns to balance the contributions of these three component networks. However, although $AVTENet_{sf}$ outperforms $AVTENet_{mv}$ and $AVTENet_{asf}$ on the faceswap test set (0.92 vs 0.87 and 0.86), it performs worse than $AVTENet_{mv}$ and $AVTENet_{asf}$ on other test sets.
$\mathbf{AVTENet_{ff}}$: Table X shows the results of $AVTENet_{ff}$ on eight different test sets. We can see that $AVTENet_{ff}$ performs best among four ensemble models on faceswap, fsgan, and Testset-II, and achieves near-perfect performance on faceswap-wav2lip, fsgan-wav2lip, wav2lip, and Testset-II. While other ensemble models may perform well on specific test sets, $AVTENet_{ff}$ maintains stable and reliable performance on all eight test sets. The good performance may be attributed to the additional linear layers that takes as input the embeddings of the penultimate layers of the VN, AN, and AVN networks, as it learns to balance the contributions of these three component networks. As shown in Figs. 8 and 9, the discriminative ability of the fused feature $\mathbf{E_{ff}}$ is further improved in the additional linear classification layer of $AVTENet_{ff}$. In summary, feature fusion is the best fusion strategy in this study.
@@
Testset-I Real 0.96 0.97 0.96 0.96
Fake 0.97 0.96 0.96
Testset-II Real 1.00 0.97 0.99 0.99
Fake 0.97 1.00 0.99
@@
Figure 8: t-SNE visualization of the fused feature $\mathbf{E_{ff}}$ in $AVTENet_{ff}$.
@@
Figure 9: t-SNE visualization of features in the linear classifier of $AVTENet_{ff}$.
IV-D5 Comparison of The Above Models
Fig. 10 depicts the accuracy of the above seven models on the two main test sets, Testset-I and Testset-II. The video-only model (VN) is good at identifying visual manipulations in videos. Likewise, the audio-only model (AN) is good at identifying manipulated acoustic content in videos. But they only focus on a single modality. In contrast to VN and AN, the AVN and all variants of AVTENet can identify acoustic and visual manipulations in videos. Although $AVTENet_{mv}$ and $AVTENet_{asf}$ have higher accuracy than $AVTENet_{ff}$ on Testset-II, the latter shows higher stability in detecting various manipulation types in videos.
Comparing Tables VII-IX with Table IV, we can see that among the variants of AVTENet, only $AVTENet_{ff}$ can outperform the VN model on all test sets, while the remaining AVTENet models perform worse than the VN model on the faceswap test set (fake samples containing only visual manipulation), despite outperforming the VN model on the other test sets. This is because the VN model only focuses on the visual modality, while the AVTENet models may be interfered with by the acoustic feature extracted from the real auditory modality when detecting visual manipulation. Furthermore, comparing Tables VII-IX with Table V, all AVTENet models outperform the AN model on all test sets except RTVC (fake samples containing only acoustic manipulation). This happens because the AN model only focuses on the acoustic modality, while the AVTENet models may be interfered with by the visual feature extracted from the real visual modality when detecting acoustic manipulation.
TABLE XII: Performance comparison of our Ensemble model and several existing unimodal, multimodal, fusion, and ensemble models. DFD in the first column refers to the deepfake detection method. “V”, “A” and “AV” stand for visual, audio and audio-visual modalities, respectively.
@@ IV-D5 Comparison of The Above Models
Unimodal[69] VGG16 V Real 0.69 0.90 0.78 0.81
Fake 0.87 0.78 0.82
Unimodal[69] Xception A Real 0.88 0.61 0.72 0.76
Fake 0.70 0.91 0.80
@@
Ensemble (Soft-Voting)[69] VGG16 AV Real 0.69 0.90 0.78 0.78
Fake 0.90 0.69 0.78
Ensemble (Hard-Voting)[69] VGG16 AV Real 0.69 0.90 0.78 0.78
Fake 0.90 0.69 0.78
Multimodal-1[69] Multimodal-1 AV Real 0.00 0.00 0.00 0.50
Fake 0.50 1.00 0.66
Multimodal-2[69] Multimodal-2 AV Real 0.71 0.59 0.64 0.67
Fake 0.65 0.76 0.70
Multimodal-3[69] CDCN AV Real 0.50 0.07 0.12 0.52
Fake 0.50 0.94 0.65
@@
Modality Mixing [70] Multimodaltrace AV Real - - - 0.93
@@
Ensemble[71] AVFakeNet AV Real - - - 0.93
@@
Fusion[72] AVoiD-DF AV Real - - - 0.84
@@
Fusion[73] MIS-AVioDD AV Real - - - 0.96
@@
IV-D6 Comparison of AVTENet with Different AVNs